OpenStudy (anonymous):

Topic: Partial Derivatives A computer company has a monthly advertising budget of $ 60,000.Its marketing department estimates that if x dollars are spent each month on advertising in newspapers and y dollars per month on television advertising, then the monthly sales is given by dollars. If the profit is 15% of sales, less the advertising cost, determine how to allocate the advertising budget in order to maximize the monthly profit

OpenStudy (amistre64):

P = S(.15) - Ac ... got that much of it :)

OpenStudy (amistre64):

x + y = 60,000

OpenStudy (amistre64):

P = S(.15) - 60,000 .... does this look right to you?

OpenStudy (anonymous):

yes i had up to there, I guess I am on the right track

OpenStudy (anonymous):

thank you very much

OpenStudy (amistre64):

is all of the 60,000 to be spent? or some portion of it on x and y?

OpenStudy (amistre64):

it doesnt say how much sales are generated by newspaper, and how much by tv.... thats whats getting me...

OpenStudy (anonymous):

no it does not so, i am assuming, that , X-Y and Y=X

OpenStudy (amistre64):

maybe have to do: y = 60,000 -x and x = 60000 -y in place of that 60,000; but that still dont seem right to me....

OpenStudy (amistre64):

your smarter at this than I am :)

OpenStudy (amistre64):

you sure there aint no pertinent information being left out?

OpenStudy (amistre64):

without knowing the sales generated by newspaper and tv, I got nothing....

OpenStudy (anonymous):

S = 90x ^1/4 y^3/4 are the constraints of x +y =60,000 F( x,y, *) = f( x,y) - * ( g(x,y)- c)

OpenStudy (anonymous):

where * is lamda

OpenStudy (amistre64):

whats your gut tell you about what to do with: S= 90 x^1/4 y^3/4 ?? Can we derive this with respect to "S" ? And dont trust my judgement on this, cause its a little above me at the moment :)

OpenStudy (amistre64):

Do we make one variable constant?

OpenStudy (anonymous):

where * is lamda